A train running along a straight track at 30m/s is slowex unformly to a stop in 44s. Find the acceleration and the stopping distance
1
Expert's answer
2020-02-05T12:27:18-0500

The acceleration:


"a=\frac{0-v}{t}=\frac{0-30}{44}=-0.68\frac{m}{s^2}"

The stopping distance:


"d=0.5vt=0.5(30)(44)=660\ m"